# BEVCar: Camera-Radar Fusion for BEV Map and Object Segmentation

### v1

@@ II Related Work
Recent advancements include full-surround view BEV perception approaches, such as CVT [20] that uses a cross-view transformer with learned positional embeddings to avoid explicit geometric modeling. In contrast, BEVFormer [4] and BEVSegFormer [21] model geometry explicitly using camera calibration parameters and propose a deformable attention-based [22] spatial cross-attention module for view unprojection. BEVFormer [4] additionally employs a temporal attention module for aggregating BEV maps over time using vehicle ego-motion, which represents the state of the art in 3D object detection. Temporal aggregation is also employed in BEVerse [11], which extends existing approaches with a motion prediction head and demonstrates that the proposed multi-task network outperforms single-task networks indicating a positive transfer among the tasks. The aforementioned approaches are often combined with novel data augmentation techniques [23], which address the view discrepancy between the image and BEV by maintaining spatial consistency across each intermediate embedding. Finally, SkyEye [24] proposes a less-constrained method that learns semantic BEV maps from labeled frontal view images by reconstructing semantic images over time. Our work leverages recent progress in monocular BEV perception and takes advantage of the radar modality for a more geometrically feasible view projection. This is achieved through a novel attention-based image lifting scheme using radar queries. Additionally, we propose to exploit existing image backbones that are pre-trained with contrastive learning to further regularize the modality-specific branches.
@@
In segmentation, initial works investigated semantic segmentation of radar point clouds [29] without complementary vision input. More recently, research towards multi-modal BEV map and object segmentation has received growing attention. The authors of the pioneering work FISHING Net [12] propose an MLP-based lifting strategy for camera features. To combine these features with radar data, which are encoded by a UNet-like network, FISHING Net performs class-based priority pooling. In contrast, Simple-BEV [6] processes the raw radar data in a rasterized BEV format and concatenates these with image features that are lifted via bilinear sampling. Although Simple-BEV targets object-agnostic vehicle segmentation, the training relies on additional instance information. Since purely concatenation-based fusion might suffer from spatial misalignment, CRN [9] employs deformable attention [22] to aggregate image and radar features. However, the method uses LSS [10] for lifting the image features and requires LiDAR during training to supervise the depth distribution network. Finally, BEVGuide [8] does not exploit further knowledge other than available during deployment. Using homography-based projection, features from the EfficientNet [30] image backbone are transformed into a scale-ambiguous top-down representation. The radar data is converted to BEV space and then encoded by two convolutional layers. In contrast to prior works, BEVGuide proposes a bottom-up lifting approach by querying the sensor features from a unified BEV space to obtain sensor-specific embeddings that are then concatenated. In this work, we further advance these ideas and utilize a more refined radar encoder that is inspired by LiDAR processing [26]. Moreover, we propose a novel lifting scheme that explicitly leverages radar points as a strong prior.
@@ III-C Sensor Fusion
Importantly, the lifted image features only serve as keys and values during the cross-attention step. In total, we utilize a cascade of six transformer blocks. Finally, we feed the output of the last block through a ResNet-18 [31] bottleneck to jointly encode the features of both modalities.
@@
We compare BEVCar with both camera-only (C) and camera-radar (C+R) BEV segmentation methods on the nuScenes [1] validation split. Simple-BEV++ is a customized Simple-BEV [6] without instance-aware training but with the same radar metadata and map segmentation head as our method. Note that CRN [9] uses LiDAR during training. The “map” metric averages the IoU of all nuScenes map classes. Previous works report predictions for fewer classes, indicated by “–”. To compare BEVCar with these methods, we provide the mean of “vehicle” and “drivable area” classes as “mIoU”.
III-D BEV Segmentation Head
We employ a single head for multi-class BEV segmentation. In detail, we utilize two convolutional layers with ReLU activations followed by a final convolutional layer to output one object class and map classes. Given the BEV space resolution, the segmentation head produces an output of size . Thus, a pixel can not only capture both a vehicle and a map class prediction but can also be assigned to multiple map categories.
Object Segmentation: For segmenting objects, we consider all vehicle-like entities, e.g., passenger cars and trucks. Unlike prior works [6], we emphasize that object-agnostic segmentation should not rely on instance-aware information during training time as this renders the application of a method less flexible. Therefore, we supervise the object channel of the segmentation head solely via the binary cross-entropy loss:
@@ III-D BEV Segmentation Head
with tunable parameter to address foreground/background imbalance.
@@ IV-A Experimental Settings
Dataset and Metrics: We evaluate our BEVCar approach on the nuScenes dataset [1] for automated urban driving in Singapore and Boston, MA, being the only publicly available dataset that provides the required sensor data and ground truth map annotations. The nuScenes dataset comprises surround-view vision from six RGB cameras and five automotive radars and provides BEV map information. For training and evaluation, we use the official training/validation split, containing 28,130 and 6,019 samples, respectively. We further categorize the validation scenes into day (4,449 samples), rain (968 samples), and night (602 samples) scenes and include this split in our code release. For object segmentation, we combine all subclasses of the “vehicle” category. For map segmentation, we consider all available classes, i.e., “drivable area”, “carpark area”, “pedestrian crossing”, “walkway”, “stop line”, “road divider”, and “lane divider”. We report individual intersection over union (IoU) metric [36] for those classes that have been addressed by prior works and refer to the mean IoU of all map classes by “map”. To compare BEVCar with previous baselines that predict fewer classes, we report the average of “vehicle” and “drivable area” as “mIoU”.
@@ IV-B Quantitative Results
We compare BEVCar to various baseline works in Tab. I. In particular, we report the performance for the camera-radar fusion methods Simple-BEV [6], BEVGuide [8], and CRN [9], which leverages depth from LiDAR during training. At the time of submission, only the authors of Simple-BEV released their code. We utilize this for an extended version Simple-BEV++ by adding the BEV map segmentation task, removing additional radar metadata (see Sec. III-A), and disregarding the instance-aware losses (see Sec. III-D). To demonstrate the advantage of radar measurements, we further compare BEVCar to the vision-only baselines CVT [20], BEVFormer [4], and variations of both Simple-BEV [6] and our proposed BEVCar.
@@ IV-C Ablations and Analysis
To further analyze our proposed BEVCar approach, we provide ablations for its components and evaluate the performance improvements over the baselines under challenging conditions.
@@
Distance-Based Object Segmentation: In Tab. III, we analyze the vehicle segmentation quality of BEVCar, its camera-only variant, Simple-BEV [6], and Simple-BEV++ for three different range intervals including -, -, and -. Note that the overall performance of Simple-BEV is slightly lower than reported in Tab. I due to rerunning the authors’ code to enable the range-based evaluation. Similar to the previous experiment, we observe that the results of the camera-only baseline significantly differ between the evaluation criteria. While the IoU in the - range is comparable to Simple-BEV, for the - range it achieves only half of the initial performance. Although the general trend is similar among all camera-radar methods, the effect is the least severe for BEVCar. To conclude, our experiment demonstrates the advantage of utilizing radar measurements to maintain object segmentation performance also at larger distances.
@@
Robustness to Weather and Illumination: Besides providing complementary information, i.e., dense RGB data versus sparse distance and velocity measurements, a core difference between cameras and radars is the source of energy utilized by the respective sensor. While passive sensors such as cameras rely on an external source like the sun, active sensors such as radars provide their own energy. Therefore, passive sensors suffer from challenging illumination conditions, e.g., faced during rain or at night. We thus emphasize that evaluating automotive perception systems specifically in these situations is imperative to fully understand their performance.
In Tab. IV and Fig. 5, we separate the previously reported metrics for BEVCar and the same baselines as in the study on the perception range into day, rain, and night. We observe that the vehicle segmentation IoU of the camera-only baseline is subject to substantial degradation during the night. In contrast, all camera-radar methods can maintain their performance, whereas BEVCar achieves the highest performance. On the other hand, the map segmentation mIoU decreases during rain and even further at night, which holds for all investigated methods. The results indicate that radar is most beneficial for object detection and less important for BEV mapping, which is expected as depth information is less important for learning a mapping of the planar map classes from the 2D image space to the BEV space than mapping objects with defined height, width, and depth parameters.
@@ V Conclusion
In this work, we introduced BEVCar addressing camera-radar fusion for BEV map and object segmentation. BEVCar comprises a new learning-based radar point encoding and leverages radar information early during the lifting step of the vision features from the image plane to the BEV space. We demonstrated that BEVCar outperforms previous camera-radar approaches when considering map and object segmentation in a joint manner. We extensively evaluated the performance in challenging weather and illumination conditions and analyzed the robustness for various perception ranges. Our results clearly demonstrate the benefit of utilizing automotive radar in addition to surround-view vision. To facilitate further research in this direction, we include our day/rain/night split of the nuScenes [1] validation data in the public release of our code. In the future, we will address robustness in case of partial or complete sensor failure, e.g., by leveraging cross-modality distillation during the training process.

### v2

@@ II Related Work
Recent advancements include full-surround view BEV perception approaches, such as CVT [20] that uses a cross-view transformer with learned positional embeddings to avoid explicit geometric modeling and exploiting this BEV representation for policy learning [21]. In contrast, BEVFormer [4] and BEVSegFormer [22] model geometry explicitly using camera calibration parameters and propose a deformable attention-based [23] spatial cross-attention module for view unprojection. BEVFormer [4] additionally employs a temporal attention module for aggregating BEV maps over time using vehicle ego-motion, which represents the state of the art in 3D object detection. Temporal aggregation is also employed in BEVerse [11], which extends existing approaches with a motion prediction head and demonstrates that the proposed multi-task network outperforms single-task networks indicating a positive transfer among the tasks. The aforementioned approaches are often combined with novel data augmentation techniques [24], which address the view discrepancy between the image and BEV by maintaining spatial consistency across each intermediate embedding. Finally, SkyEye [25] proposes a less-constrained method that learns semantic BEV maps from labeled frontal view images by reconstructing semantic images over time. Our work leverages recent progress in monocular BEV perception and takes advantage of the radar modality for a more geometrically feasible view projection. This is achieved through a novel attention-based image lifting scheme using radar queries. Additionally, we propose to exploit existing image backbones that are pre-trained with contrastive learning to further regularize the modality-specific branches.
@@
In segmentation, initial works investigated semantic segmentation of radar point clouds [30] without complementary vision input. More recently, research towards multi-modal BEV map and object segmentation has received growing attention. The authors of the pioneering work FISHING Net [12] propose an MLP-based lifting strategy for camera features. To combine these features with radar data, which are encoded by a UNet-like network, FISHING Net performs class-based priority pooling. In contrast, Simple-BEV [6] processes the raw radar data in a rasterized BEV format and concatenates these with image features that are lifted via bilinear sampling. Although Simple-BEV targets object-agnostic vehicle segmentation, the training relies on additional instance information for object center and pixel offset prediction. Since purely concatenation-based fusion might suffer from spatial misalignment, CRN [9] employs deformable attention [23] to aggregate image and radar features. However, the method uses LSS [10] for lifting the image features and requires LiDAR during training to supervise the depth distribution network. Finally, BEVGuide [8] does not exploit further knowledge other than available during deployment. Using homography-based projection, features from the EfficientNet [31] image backbone are transformed into a scale-ambiguous top-down representation. The radar data is converted to BEV space and then encoded by two convolutional layers. In contrast to prior works, BEVGuide proposes a bottom-up lifting approach by querying the sensor features from a unified BEV space to obtain sensor-specific embeddings that are then concatenated. In this work, we further advance these ideas and utilize a more refined radar encoder that is inspired by LiDAR processing [27]. Moreover, we propose a novel lifting scheme that explicitly leverages radar points as a strong prior.
@@ III-C Sensor Fusion
Importantly, the lifted image features only serve as keys and values during the cross-attention step. In total, we utilize a cascade of six transformer blocks. Finally, to encode the features of both modalities in a joint manner, we feed the output of the last block through a ResNet-18 [32] bottleneck, referred to as BEV encoder in Fig. 2.
@@
We compare BEVCar with both camera-only (C) and camera-radar (C+R) BEV segmentation methods on the nuScenes [2] validation split. Simple-BEV++ is a customized Simple-BEV [6] without instance-aware training but with the same radar metadata and map segmentation head as our method. Note that CRN [9] uses LiDAR during training. The “map” metric averages the IoU of all nuScenes map classes. Previous works report predictions for fewer classes, indicated by “–”. To compare BEVCar with these methods, we provide the mean of “vehicle” and “drivable area” classes as “mIoU”. Bold and underlined values denote the best and second-best metrics per column, respectively.
III-D Segmentation Head
We employ a single head for multi-class BEV semantic segmentation. In detail, we utilize two convolutional layers with ReLU activations followed by a final convolutional layer to output one object class and map classes. Given the BEV space resolution, the segmentation head produces an output of size . Thus, a pixel can not only capture both a vehicle and a map class prediction but can also be assigned to multiple map categories.
Object Segmentation: For segmenting objects, we consider all vehicle-like entities, e.g., passenger cars and trucks. Unlike prior works [6], we emphasize that object-agnostic segmentation should not rely on instance-aware information during training time as this renders the application of a method less flexible due to requiring additional annotations. Therefore, we supervise the object channel of the segmentation head solely via the binary cross-entropy loss:
@@ III-D Segmentation Head
with tunable parameter to address the foreground-background imbalance.
@@ IV-A Experimental Settings
Dataset and Metrics: We evaluate our BEVCar approach on the nuScenes dataset [2] for automated urban driving in Singapore and Boston, MA, being the only publicly available dataset that provides the required sensor data and ground truth map annotations. The nuScenes dataset comprises surround-view vision from six RGB cameras and five automotive radars and provides BEV map information. For training and evaluation, we use the official training/validation split, containing 28,130 and 6,019 samples, respectively. We further categorize the validation scenes into day (4,449 samples), rain (968 samples), and night (602 samples) scenes and include this split in our code release. For object segmentation, we combine all subclasses of the “vehicle” category. For map segmentation, we consider all available classes, i.e., “drivable area”, “carpark area”, “pedestrian crossing”, “walkway”, “stop line”, “road divider”, and “lane divider”. We report individual intersection over union (IoU) metrics [37] for those classes that have been addressed by prior works and refer to the mean IoU of all map classes by “map”. To compare BEVCar with previous baselines that predict fewer classes, we report the average of “vehicle” and “drivable area” as “mIoU”.
@@ IV-B Quantitative Results
We compare BEVCar to various baseline works in Tab. I, including the camera-radar fusion methods Simple-BEV [6], BEVGuide [8], and CRN [9], which leverages depth from LiDAR during training. At the time of submission, only the authors of Simple-BEV released their code. We utilize this for an extended version Simple-BEV++ by adding the BEV map segmentation task, removing additional radar metadata (see Sec. III-A), and disregarding the instance-aware losses (see Sec. III-D). To demonstrate the advantage of radar measurements, we further compare BEVCar to the vision-only baselines CVT [20], BEVFormer [4], and variations of both Simple-BEV [6] and our proposed BEVCar.
@@ IV-C Ablations and Analysis
To further analyze our proposed BEVCar approach, we provide ablations for its components and evaluate its performance under challenging conditions.
@@
Distance-Based Object Segmentation: In Tab. III, we analyze the vehicle segmentation quality of BEVCar, its camera-only variant, Simple-BEV [6], and Simple-BEV++ for three different range intervals including -, -, and -. Note that the overall performance of Simple-BEV is slightly lower than reported in Tab. I due to rerunning the authors’ code to enable the range-based evaluation. Generally, we observe that the results of the camera-only baseline significantly differ between the evaluation criteria. While the IoU in the - range is comparable to Simple-BEV, for the - range it achieves only half of the initial performance. Although the general trend is similar among all camera-radar methods, the effect is the least severe for BEVCar. Our experiment demonstrates the advantage of utilizing radar measurements to maintain object segmentation performance also at larger distances.
@@
| Image | Deformable | Runtime | ||
| Method | Backbone | Attention | ms | FPS |
| Simple-BEV [6] | ResNet-101 | 18 | 54.9 | |
| BEVCar | ResNet-101 | ✓ | 137 | 7.3 |
| BEVCar (camera) | ViT-B/14 | ✓ | 352 | 2.8 |
| Simple-BEV++ | ViT-B/14 | ✓ | 277 | 3.6 |
| BEVCar (ours) | ViT-B/14 | ✓ | 382 | 2.6 |
Runtime of a forward pass measured on an Nvidia A100 GPU.
Robustness to Weather and Illumination: Besides providing complementary information, i.e., dense RGB data versus sparse distance and velocity measurements, a core difference between cameras and radars is the source of energy utilized by the respective sensor. While passive sensors such as cameras rely on an external source like the sun, active sensors such as radars provide their own energy. Therefore, passive sensors suffer from challenging illumination conditions, e.g., faced during rain or at night. We thus emphasize that evaluating automotive perception systems specifically in these situations is imperative to understand their performance fully.
In Tab. IV, Fig. 1, and Fig. 5, we separate the previously reported metrics for BEVCar and the same baselines as in the study on the perception range into day, rain, and night. We observe that the vehicle segmentation IoU of the camera-only baseline is subject to substantial degradation during the night. In contrast, all camera-radar methods can maintain their performance, whereas BEVCar achieves the highest performance. On the other hand, the map segmentation mIoU decreases during rain and even further at night, which holds for all investigated methods. The results indicate that radar is most beneficial for object detection and less relevant for BEV mapping, which is expected as depth information is less important for learning a mapping of the planar map classes from the 2D image space to the BEV space than mapping objects with defined height, width, and depth parameters.
Runtime Analysis: In Tab. V, we report the runtime of our proposed BEVCar and multiple baseline methods, measured on an Nvidia A100 GPU and averaged over the validation split. In contrast to Harley et al. [6], we only consider the forward pass without data loading and loss calculation. Most notable is the slow-down caused by deformable attention [23] employed in our proposed lifting module as well as in the ViT adapter. Importantly, in comparison to the vision-only baseline, including the radar information does not result in a significantly higher runtime. Note that the frequency of the synchronized keyframes in the nuScenes [2] dataset is .
@@ V Conclusion
In this work, we introduced BEVCar addressing camera-radar fusion for BEV map and object segmentation. BEVCar comprises a new learning-based radar point encoding and leverages radar information early during the lifting step of the vision features from the image plane to the BEV space. We demonstrated that BEVCar outperforms previous camera-radar approaches when jointly considering map and object segmentation. We extensively evaluated the performance in challenging weather and illumination conditions and analyzed the robustness for various perception ranges. Our results clearly demonstrate the benefit of utilizing automotive radar in addition to surround-view vision. To facilitate further research in this direction, we include our day/rain/night split of the nuScenes [2] validation data in the public release of our code. In the future, we will address robustness in case of partial or complete sensor failure, e.g., by leveraging cross-modality distillation during training of the network.